OpenStudy (anonymous):

What are the intercepts of the equation -2x + 5y = 20?

OpenStudy (anonymous):

5y=2x+20 y=(2/5)x+4 y-intercept:y=4 x-intercept:x=-10

OpenStudy (anonymous):

A. (-10, 0), (0, 4) B. (0, -10), (4, 0) C. (10, 0), (0, -4) D. (0, 10), (-4, 0) SO IS IT A ?

OpenStudy (anonymous):

yes ,A